OpenStudy (anonymous):

Daniel traveled 300 miles in 6 hours. Using the formula d=rt where d is distance, r is rate and t is time, what was Daniel's rate for the trip?

OpenStudy (anonymous):

d=300 t=6 r=d/t=300/6=50 rate is 50